Table 24-11. Megan's salary for three consecutive years, along with other values, are presented in the table below.

Year
2011 2012 2013
Salary $65,000 $72,000 $76,000
Consumer Price Index 226 230 235
Real Interest Rate 2.5 percent 2.7 percent 1.8 percent

Refer to Table 24-11. Suppose the consumer price index for 2013 is not necessarily 235. If the nominal interest rate for 2013 is 7.3 percent , then the consumer price index for 2013 is, in fact,
Answer #1

Nominal interest rate for 2013 = 7.3%

Real interest rate for 2013 = 1.8%

So the rate of inflation = Nominal - Real = 7.3-1.8 = 5.5%

So the CPI should also increase by 5.5% in 2013.

So the CPI in 2013 = 230 * 1.055 = 242.65